OpenStudy (mathmath333):

If three dice are rolled .what is the number of possible outcomes in which exactly one dice shows the number of 2 ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{If three dice are rolled .what is the number of possible outcomes}\hspace{.33em}\\~\\ & \normalsize \text{ in which exactly one dice shows the number of 2 ?}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (welshfella):

hay i hate these!!

OpenStudy (mathmath333):

u hate these questions

OpenStudy (welshfella):

yea 2 n2 n2 n2 2 n2 n2 n2 2 that's 3 for a start

OpenStudy (anonymous):

"one dice" = die

OpenStudy (mathmath333):

what was that from satellite

OpenStudy (welshfella):

and there are 5 ways of getting not 2 some call a dice a die

OpenStudy (welshfella):

is it 3*5*5 - or is that too simple!!!

OpenStudy (anonymous):

i think it may be \(3\times 36\)

OpenStudy (mathmath333):

it is 75

OpenStudy (anonymous):

oh no, EXACTLY one die is a five

OpenStudy (anonymous):

is a two whatever go with @welshfella

imqwerty (imqwerty):

one die = a die ( ͡° ͜ʖ ͡°)

OpenStudy (anonymous):

first die is a 2, then 25 possible outcomes second die is a 2, then 25 third die is a 2, then 25 \[3\times 25\] or as @welshfella said \(3\times 5^2\)

OpenStudy (mathmath333):



OpenStudy (mathmath333):

5*5*3=75 got it

imqwerty (imqwerty):

u just fix one die on which u have a 2 and then find out the number of all possible combinations that can be made with the other 2 dies (exclude the 2) and then multiply the answer by 3 :P

OpenStudy (mathmath333):

i like the drawing tool

imqwerty (imqwerty):

i think ur answer is correct :)

imqwerty (imqwerty):

lol XD

OpenStudy (mathmath333):

haha